# FDDM: Unsupervised Medical Image Translation with a Frequency-Decoupled Diffusion Model

### v2

@@ Abstract
Diffusion models have demonstrated significant potential in producing high-quality images in medical image translation to aid disease diagnosis, localization, and treatment. Nevertheless, current diffusion models have limited success in achieving faithful image translations that can accurately preserve the anatomical structures of medical images, especially for unpaired datasets. The preservation of structural and anatomical details is essential to reliable medical diagnosis and treatment planning, as structural mismatches can lead to disease misidentification and treatment errors. In this study, we introduce the Frequency Decoupled Diffusion Model (FDDM) for MR-to-CT conversion. FDDM first obtains the anatomical information of the CT image from the MR image through an initial conversion module. This anatomical information then guides a subsequent diffusion model to generate high-quality CT images. Our diffusion model uses a dual-path reverse diffusion process for low-frequency and high-frequency information, achieving a better balance between image quality and anatomical accuracy. We extensively evaluated FDDM using public datasets for brain MR-to-CT and pelvis MR-to-CT translations, demonstrating its superior performance to other GAN-based, VAE-based, and diffusion-based models. The evaluation metrics included Fréchet Inception Distance (FID), Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index Measure (SSIM). FDDM achieved the best scores on all metrics for both datasets, particularly excelling in FID, with scores of 25.9 for brain data and 29.2 for pelvis data, significantly outperforming other methods. These results demonstrate that FDDM can generate high-quality target domain images while maintaining the accuracy of translated anatomical structures.
{IEEEkeywords}
Medical Image Translation, Diffusion Model, Generative Model
@@
1 Related Work
1.1 Unsupervised Medical Image Translation.
@@ 1 Related Work
1.2 Diffusion Models for Image Translation
@@ 2 Method
2.1 Initial Conversion Module
@@
$\begin{split}\mathcal{L}_{\text{\tiny VAE}_{A}}=&\lambda_{1}\text{KL}(q_{A}(L_{A}|A)||\mathcal{N}(0,I))\\ &-\lambda_{2}\mathbb{E}_{L_{A}\sim q_{A}(L_{A}|A)}[|A-D_{A}(L_{A})|],\end{split}$ (8)
$\begin{split}\mathcal{L}_{\text{\tiny VAE}_{B}}=&\lambda_{1}\text{KL}(q_{B}(L_{B}|B)||\mathcal{N}(0,I))\\ &-\lambda_{2}\mathbb{E}_{L_{B}\sim q_{B}(L_{B}|B)}[|B-D_{B}(L_{B})|],\end{split}$ (9)
@@
$\begin{split}\mathcal{L}_{\text{\tiny CC}_{A}}&=\lambda_{1}\text{KL}(q_{A}(L_{A}|A)||\mathcal{N}(0,I))\\ &+\lambda_{1}\text{KL}(q_{B}(L_{B}|D_{B}(E_{A}(A))))||\mathcal{N}(0,I))\\ &-\lambda_{2}\mathbb{E}_{L_{B}\sim q_{B}(L_{B}|D_{B}(E_{A}(A)))}[|A-D_{A}(L_{B})|],\end{split}$ (12)
$\begin{split}\mathcal{L}_{\text{\tiny CC}_{B}}&=\lambda_{1}\text{KL}(q_{B}(L_{B}|B)||\mathcal{N}(0,I))\\ &+\lambda_{1}\text{KL}(q_{A}(L_{A}|D_{A}(E_{B}(B))))||\mathcal{N}(0,I))\\ &-\lambda_{2}\mathbb{E}_{L_{A}\sim q_{A}(L_{A}|D_{A}(E_{B}(B)))}[|B-D_{B}(L_{A})|].\end{split}$ (13)
@@
$\begin{split}\mathcal{L}_{\text{\tiny RC}_{A}}=&-\lambda_{2}\mathbb{E}[|LR(D_{B}(E_{A}(RR(A))))-D_{B}(E_{A}(A))|]\\ &-\lambda_{2}\mathbb{E}[|LR(D_{B}(RR(E_{A}(A))))-D_{B}(E_{A}(A))|],\end{split}$ (14)
$\begin{split}\mathcal{L}_{\text{\tiny RC}_{B}}=&-\lambda_{2}\mathbb{E}[|LR(D_{A}(E_{B}(RR(B))))-D_{A}(E_{B}(B))|]\\ &-\lambda_{2}\mathbb{E}[|LR(D_{A}(RR(E_{B}(B))))-D_{A}(E_{B}(B))|].\end{split}$ (15)
@@
Subsequently, we perform a forward diffusion (i.e., low-pass filtering) to convert the coarse CT $\overline{Y}$ to the noisy CT $y_{T_{s}}$ as the starting point for reverse diffusion. As shown in Fig. 1, $S(\overline{Y})$ represents the local structure, and $y_{T_{s}}$ represents the overall CT anatomy pattern. Together, they serve as the necessary anatomical information of the CT to guide the reverse diffusion process. Even though we already have a coarse CT $\overline{Y}$, the following diffusion process is still necessary for further image correction and enhancement, since the image quality of the coarse CT $\overline{Y}$ from the VAE-based initial conversion module is low with data distribution mismatches. The related experimental validation is presented in Section 3.3.
2.2 Forward Diffusion Process
@@
Knowing that the forward diffusion process approximates a low-pass filter, a to-be-determined parameter is the threshold for the low-pass filter, i.e., how many steps the forward diffusion should take. Since $\alpha_{t}$ is a decreasing sequence, $\text{SNR}(y_{t},f)$ decreases as the forward diffusion step $t$ increases. Therefore, if the number of steps is too large, very little image information would be retained to effectively guide the reverse diffusion process. On the other hand, if the number of steps is too small, it won’t adequately disrupt/override the data distribution errors in the coarse CT $\overline{Y}$ to allow further image correction and enhancement. We hypothesize that there is an optimal number of steps $T_{s}$, which will be determined through experiments (as discussed in Section 3.4).
2.3 Dual-Path Reverse Diffusion Process
@@
Table 1: Comparison of different methods on the brain dataset. Downward arrows indicate ”lower is better” and upward arrows indicate ”higher is better”. The results for FDDM are obtained based on $T_{s}=300$.
Method FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
CycleGAN 62.4051 0.8788 $\pm$ 0.0446 36.5048 $\pm$ 2.3845
GcGAN 60.0342 0.8841 $\pm$ 0.0462 36.9609 $\pm$ 2.6264
RegGAN 73.7575 0.8187 $\pm$ 0.0492 35.7266 $\pm$ 1.2574
UNIT 49.7074 0.8960 $\pm$ 0.0427 36.8695 $\pm$ 2.6763
MUNIT 72.8495 0.8449 $\pm$ 0.0596 35.7886 $\pm$ 1.9414
SDEdit 75.9913 0.7385 $\pm$ 0.0978 35.3694 $\pm$ 2.7448
SynDiff 70.5355 0.8632 $\pm$ 0.0484 37.1294 $\pm$ 2.8835
FDDM 25.8565 0.9144 $\pm$ 0.0379 38.0815 $\pm$ 2.8457
@@
3 Experiments
3.1 Experimental and Dataset Setup
In our study, we used the SynthRAD2023 dataset [^1] https://synthrad2023.grand-challenge.org/, which includes 180 pairs of 3D scans from the brain (MR and CT) and 180 pairs of 3D scans from the pelvis (MR and CT). For our experimental setup, we randomly split these images, reserving 10 pairs of CT and MR scans each from the brain and pelvis as our test set, and 10 pairs each for validation (per needs of comparison methods). The remaining 160 paired MR-CT scans were used for training. Each 3D image was processed into 2D slices. To simulate unpaired (unsupervised) training, we shuffled the slices and removed the pairing information from the training dataset. All slices were uniformly resized to a resolution of 256 × 256 pixels. For MR images, we applied intensity clipping, removing the bottom 0.5% of intensity values. For CT images, the Hounsfield Unit (HU) values were clipped to the range of [-1000, 1000]. All images were then rescaled to [-1, 1]. The training MR and CT slices were used to train the initial conversion module, while only the training CT slices were used to train the diffusion model, which does not need any MR input. In our implementation, several parameters were determined based on default values from previous methods. Our diffusion model has $T=1000$ steps. The learning rates for the VAE-based initial conversion module and the diffusion model were set to $1\times 10^{-4}$ and $2\times 10^{-4}$, respectively. Exponential Moving Average (EMA) was applied to the model parameters with a decay factor of 0.9999. The loss weights for the initial conversion module followed the widely used VAE framework, with $\lambda_{1}$ set to 0.01 and $\lambda_{2}$ set to 10. The value of $T_{s}$ during the inference process was set to 500, based on parameter studies (Section III). We used the Pytorch library to implement the FDDM framework. Experiments were conducted on an NVIDIA A100 80G GPU card. We compared FDDM with other state-of-the-art methods (GcGAN, RegGAN, CycleGAN, UNIT, MUNIT, SynDiff, SDEdit) using official open-source code[22, 23, 25, 26, 27].
@@ 3.1 Experimental and Dataset Setup
Table 2: Comparison of different methods on the pelvis MR-to-CT translation dataset. Downward arrows indicate ”lower is better” and upward arrows indicate ”higher is better”. The results for FDDM are obtained based on $T_{s}=300$.
Method FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
CycleGAN 77.6318 0.8898 $\pm$ 0.0233 36.3827 $\pm$ 0.9034
GcGAN 43.7356 0.9091 $\pm$ 0.0231 37.3536 $\pm$ 0.9266
RegGAN 56.0113 0.8871 $\pm$ 0.0371 36.8078 $\pm$ 1.0759
UNIT 44.5953 0.9050 $\pm$ 0.0214 37.2001 $\pm$ 0.8804
MUNIT 74.3732 0.8876 $\pm$ 0.0221 36.0936 $\pm$ 0.8303
SDEdit 73.1082 0.6964 $\pm$ 0.1578 33.5062 $\pm$ 1.5040
SynDiff 74.6271 0.8842 $\pm$ 0.0236 36.7435 $\pm$ 0.9471
FDDM 29.2003 0.9184 $\pm$ 0.0210 37.6832 $\pm$ 1.0028
@@
3.2 Comparison with Other Methods
@@
3.2.1 Brain Dataset
Table 1 provides a quantitative comparison of FDDM with other image conversion methods on the brain dataset using FID, SSIM, and PSNR metrics. FDDM achieved an FID score of 25.8565, substantially outperforming other methods, demonstrating its capability to generate highly realistic images. SSIM and PSNR, which can be considered as measures of anatomical accuracy, also showed that FDDM achieved the best performance, consistent with the visual comparison results in Fig. 4. Both the quantitative and visual results indicated that SDEdit performed the worst. SDEdit uses a diffusion model trained on CT images and directly performs forward and reverse diffusion on the source MR image. Essentially, SDEdit uses the low-frequency information extracted from the source MR image (as the forward diffusion process approximates a low-pass filter) to infer the CT image, although its inference capability is gained from training with low-frequency CT image information. However, the low-frequency information between CT and MR images differs significantly, resulting in substantially degraded results from SDEdit. Another diffusion model-based method, SynDiff, showed moderate performance because it uses CycleGAN to synthesize paired MR-CT data to avoid the challenges of unsupervised training, gaining a relative advantage in structural preservation. However, introducing CycleGAN for paired CT synthesis imposed an upper limit on the imaging realism, which aligns with our experimental results showing that SynDiff’s performance was close to but slightly worse than CycleGAN. The VAE-based UNIT achieved a performance second only to our model, supporting the rationale of using VAE as the backbone structure for our initial conversion module and applying the diffusion model to further correct the residual errors.
@@ 3.2.1 Brain Dataset
Step FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$
without RC 48.6921 0.9014 $\pm$ 0.0435 37.2939 $\pm$ 2.5128
with RC 39.8187 0.9123 $\pm$ 0.0386 37.7208 $\pm$ 2.7626

### v3

@@ Abstract
Diffusion models have demonstrated significant potential in producing high-quality images in medical image translation to aid disease diagnosis, localization, and treatment. Nevertheless, current diffusion models have limited success in achieving faithful image translations that can accurately preserve the anatomical structures of medical images, especially for unpaired datasets. The preservation of structural and anatomical details is essential to reliable medical diagnosis and treatment planning, as structural mismatches can lead to disease misidentification and treatment errors. In this study, we introduce the Frequency Decoupled Diffusion Model (FDDM) for MR-to-CT conversion. FDDM first obtains the anatomical information of the CT image from the MR image through an initial conversion module. This anatomical information then guides a subsequent diffusion model to generate high-quality CT images. Our diffusion model uses a dual-path reverse diffusion process for low-frequency and high-frequency information, achieving a better balance between image quality and anatomical accuracy. We extensively evaluated FDDM using public datasets for brain MR-to-CT and pelvis MR-to-CT translations, demonstrating its superior performance to other GAN-based, VAE-based, and diffusion-based models. The evaluation metrics included Fréchet Inception Distance (FID), mean absolute error (MAE), mean squared error (MSE), and Structural Similarity Index Measure (SSIM). FDDM achieved the best scores on all metrics for both datasets, particularly excelling in FID, with scores of 25.9 for brain data and 29.2 for pelvis data, significantly outperforming other methods. These results demonstrate that FDDM can generate high-quality target domain images while maintaining the accuracy of translated anatomical structures.
Keywords: Medical Image Translation, Diffusion Model, Generative Model
I. Introduction
@@
II. Related Work
II.A. Unsupervised Medical Image Translation.
@@ II. Related Work
II.B. Diffusion Models for Image Translation
@@ III. Method
III.A. Initial Conversion Module
@@
$\begin{split}\mathcal{L}_{\text{\tiny VAE}_{A}}=&\lambda_{1}\text{KL}(q_{A}(L_{A}|A)||\mathcal{N}(0,I))\\ &+\lambda_{2}\mathbb{E}_{L_{A}\sim q_{A}(L_{A}|A)}[|A-D_{A}(L_{A})|],\end{split}$ (8)
$\begin{split}\mathcal{L}_{\text{\tiny VAE}_{B}}=&\lambda_{1}\text{KL}(q_{B}(L_{B}|B)||\mathcal{N}(0,I))\\ &+\lambda_{2}\mathbb{E}_{L_{B}\sim q_{B}(L_{B}|B)}[|B-D_{B}(L_{B})|],\end{split}$ (9)
@@
$\begin{split}\mathcal{L}_{\text{\tiny CC}_{A}}&=\lambda_{1}\text{KL}(q_{A}(L_{A}|A)||\mathcal{N}(0,I))\\ &+\lambda_{1}\text{KL}(q_{B}(L_{B}|D_{B}(E_{A}(A))))||\mathcal{N}(0,I))\\ &+\lambda_{2}\mathbb{E}_{L_{B}\sim q_{B}(L_{B}|D_{B}(E_{A}(A)))}[|A-D_{A}(L_{B})|],\end{split}$ (12)
$\begin{split}\mathcal{L}_{\text{\tiny CC}_{B}}&=\lambda_{1}\text{KL}(q_{B}(L_{B}|B)||\mathcal{N}(0,I))\\ &+\lambda_{1}\text{KL}(q_{A}(L_{A}|D_{A}(E_{B}(B))))||\mathcal{N}(0,I))\\ &+\lambda_{2}\mathbb{E}_{L_{A}\sim q_{A}(L_{A}|D_{A}(E_{B}(B)))}[|B-D_{B}(L_{A})|].\end{split}$ (13)
@@
$\begin{split}\mathcal{L}_{\text{\tiny RC}_{A}}=&\lambda_{2}\mathbb{E}[|LR(D_{B}(E_{A}(RR(A))))-D_{B}(E_{A}(A))|]\\ &+\lambda_{2}\mathbb{E}[|LR(D_{B}(RR(E_{A}(A))))-D_{B}(E_{A}(A))|],\end{split}$ (14)
$\begin{split}\mathcal{L}_{\text{\tiny RC}_{B}}=&\lambda_{2}\mathbb{E}[|LR(D_{A}(E_{B}(RR(B))))-D_{A}(E_{B}(B))|]\\ &+\lambda_{2}\mathbb{E}[|LR(D_{A}(RR(E_{B}(B))))-D_{A}(E_{B}(B))|].\end{split}$ (15)
@@
Subsequently, we perform a forward diffusion (i.e., low-pass filtering) to convert the coarse CT $\overline{Y}$ to the noisy CT $y_{T_{s}}$ as the starting point for reverse diffusion. As shown in Fig. 1, $S(\overline{Y})$ represents the local structure, and $y_{T_{s}}$ represents the overall CT anatomy pattern. Together, they serve as the necessary anatomical information of the CT to guide the reverse diffusion process. Even though we already have a coarse CT $\overline{Y}$, the following diffusion process is still necessary for further image correction and enhancement, since the image quality of the coarse CT $\overline{Y}$ from the VAE-based initial conversion module is low with data distribution mismatches. The related experimental validation is presented in Section IV.C..
III.B. Forward Diffusion Process
@@
Knowing that the forward diffusion process approximates a low-pass filter, a to-be-determined parameter is the threshold for the low-pass filter, i.e., how many steps the forward diffusion should take. Since $\alpha_{t}$ is a decreasing sequence, $\text{SNR}(y_{t},f)$ decreases as the forward diffusion step $t$ increases. Therefore, if the number of steps is too large, very little image information would be retained to effectively guide the reverse diffusion process. On the other hand, if the number of steps is too small, it won’t adequately disrupt/override the data distribution errors in the coarse CT $\overline{Y}$ to allow further image correction and enhancement. We hypothesize that there is an optimal number of steps $T_{s}$, which will be determined through experiments (as discussed in Section IV.D.).
III.C. Dual-Path Reverse Diffusion Process
@@
IV. Experiments
@@
Table 1: Comparison of different methods on the brain dataset. Downward arrows indicate ”lower is better” and upward arrows indicate ”higher is better”. The results for FDDM are obtained based on $T_{s}=300$. Values marked with * indicate statistical significance ($p<0.05$) using the Wilcoxon signed-rank test compared to FDDM.
Method FID $\downarrow$ SSIM $\uparrow$ MSE (HU) $\downarrow$ MAE (HU) $\downarrow$
CycleGAN 62.4051 0.8704* $\pm$ 0.0510 17999.873* $\pm$ 7474.353 39.261* $\pm$ 13.511
GcGAN 60.0342 0.8786* $\pm$ 0.0504 19875.950* $\pm$ 10129.532 39.842* $\pm$ 16.025
RegGAN 73.7575 0.8033* $\pm$ 0.0585 97734.234* $\pm$ 86313.220 102.163* $\pm$ 59.181
UNIT 49.7074 0.8912* $\pm$ 0.0459 17454.355* $\pm$ 8741.541 37.034* $\pm$ 14.508
MUNIT 72.8495 0.8338* $\pm$ 0.0646 30446.590* $\pm$ 14410.634 54.542* $\pm$ 20.573
SDEdit 75.9913 0.7057* $\pm$ 0.1085 106480.305* $\pm$ 59253.810 118.759* $\pm$ 52.682
SynDiff 70.5355 0.8503* $\pm$ 0.0553 42668.074* $\pm$ 21847.432 59.746* $\pm$ 23.925
FDDM 25.8565 0.9089 $\pm$ 0.0417 12345.224 $\pm$ 7052.862 28.510 $\pm$ 12.651
IV.A. Experimental and Dataset Setup
In our study, we used the SynthRAD2023 dataset [33, 34], which includes 180 pairs of 3D scans from the brain (MR and CT) and 180 pairs of 3D scans from the pelvis (MR and CT). For our experimental setup, we randomly split these images, reserving 10 pairs of CT and MR scans each from the brain and pelvis as our test set, and 10 pairs each for validation (per needs of comparison methods). The remaining 160 paired MR-CT scans were used for training. Each 3D image was processed into 2D slices. To simulate unpaired (unsupervised) training, we shuffled the slices and removed the pairing information from the training dataset. All slices were uniformly resized to a resolution of 256 × 256 pixels. For MR images, we applied intensity clipping, removing the bottom 0.5% of intensity values. For CT images, the Hounsfield Unit (HU) values were clipped to the range of [-1000, 1000]. All images were then rescaled to [-1, 1]. The training MR and CT slices were used to train the initial conversion module, while only the training CT slices were used to train the diffusion model, which does not need any MR input. In our implementation, several parameters were determined based on default values from previous methods. Our diffusion model has $T=1000$ steps. The learning rates for the VAE-based initial conversion module and the diffusion model were set to $1\times 10^{-4}$ and $2\times 10^{-4}$, respectively. Exponential Moving Average (EMA) was applied to the model parameters with a decay factor of 0.9999. The loss weights for the initial conversion module followed the widely used VAE framework, with $\lambda_{1}$ set to 0.01 and $\lambda_{2}$ set to 10. The value of $T_{s}$ during the inference process was set to 500, based on parameter studies (Section III). We used the Pytorch library to implement the FDDM framework. Experiments were conducted on an NVIDIA A100 80G GPU card. We compared FDDM with other state-of-the-art methods (GcGAN, RegGAN, CycleGAN, UNIT, MUNIT, SynDiff, SDEdit) using official open-source code [22, 23, 25, 26, 27].
Table 2: Comparison of different methods on the pelvis MR-to-CT translation dataset. Downward arrows indicate ”lower is better” and upward arrows indicate ”higher is better”. The results for FDDM are obtained based on $T_{s}=300$. Values marked with * indicate statistical significance ($p<0.05$) using the Wilcoxon signed-rank test compared to FDDM.
Method FID $\downarrow$ SSIM $\uparrow$ MSE (HU) $\downarrow$ MAE (HU) $\downarrow$
CycleGAN 77.6318 0.8840* $\pm$ 0.0256 13001.760* $\pm$ 11438.179 30.140* $\pm$ 11.539
GcGAN 43.7356 0.9014* $\pm$ 0.0257 12104.721* $\pm$ 11375.784 26.564* $\pm$ 11.674
RegGAN 56.0113 0.8763* $\pm$ 0.0412 17782.205* $\pm$ 19752.328 34.128* $\pm$ 21.893
UNIT 44.5953 0.8973* $\pm$ 0.0238 11522.925* $\pm$ 10758.339 26.997* $\pm$ 10.821
MUNIT 74.3732 0.8726* $\pm$ 0.0248 15659.264* $\pm$ 11134.156 34.739* $\pm$ 11.051
SDEdit 73.1082 0.6726* $\pm$ 0.1569 72268.590* $\pm$ 54065.650 115.527* $\pm$ 71.631
SynDiff 74.6271 0.8687* $\pm$ 0.0255 16065.846* $\pm$ 10392.632 33.364* $\pm$ 10.664
FDDM 29.2003 0.9117 $\pm$ 0.0235 8561.108 $\pm$ 10817.484 22.254 $\pm$ 10.944
@@ IV.A. Experimental and Dataset Setup
IV.B. Comparison with Other Methods
@@
IV.B.1. Brain Dataset
Table 1 provides a quantitative comparison of FDDM with other image conversion methods on the brain dataset using FID, SSIM, MSE in CT Hounsfield unit (HU), and MAE in CT Hounsfield unit (HU). FDDM achieved an FID score of 25.8565, substantially outperforming other methods, demonstrating its capability to generate highly realistic images. SSIM, MSE, and MAE, which can be considered as measures of anatomical accuracy, also showed that FDDM achieved the best performance, consistent with the visual comparison results in Fig. 4. Both the quantitative and visual results indicated that SDEdit performed the worst. SDEdit uses a diffusion model trained on CT images and directly performs forward and reverse diffusion on the source MR image. Essentially, SDEdit uses the low-frequency information extracted from the source MR image (as the forward diffusion process approximates a low-pass filter) to infer the CT image, although its inference capability is gained from training with low-frequency CT image information. However, the low-frequency information between CT and MR images differs significantly, resulting in substantially degraded results from SDEdit. Another diffusion model-based method, SynDiff, showed moderate performance because it uses CycleGAN to synthesize paired MR-CT data to avoid the challenges of unsupervised training, gaining a relative advantage in structural preservation. However, introducing CycleGAN for paired CT synthesis imposed an upper limit on the imaging realism, which aligns with our experimental results showing that SynDiff’s performance was close to but slightly worse than CycleGAN. The VAE-based UNIT achieved a performance second only to our model, supporting the rationale of using VAE as the backbone structure for our initial conversion module and applying the diffusion model to further correct the residual errors.
@@ IV.B.1. Brain Dataset
Step FID $\downarrow$ SSIM $\uparrow$ MSE (HU) $\downarrow$ MAE (HU) $\downarrow$
without RC 48.6921 0.8980 $\pm$ 0.0454 14615.918 $\pm$ 7824.331 33.0155 $\pm$ 13.7888
with RC 39.8187 0.9126 $\pm$ 0.0405 11755.669 $\pm$ 6842.154 28.2204 $\pm$ 12.5264